A glass capillary tube with a diameter of 0.43 mm and length 8 cm is filled with a salt solution with a resistivity of 0.14 Ω · m. What is the resistance?
  Ω
Answer #1

Here,

radius ,r = 0.43/2 = 0.215 mm

L = 8 cm = 0.08 m

Resistivity , p = 0.14 Ohm.m

Resistance = p * L/A

Resistance = 0.14 * 0.08/(pi * (0.215 *10^-3)^2)

Resistance = 77138 Ohm

the Resistance is 77138 Ohm